Compute the corresponding (OH-) or (H3O+) in each of the following conditions. Describe the condition as acidic or base.


1. (H3O+) = 3.7 x 10-10 M

2. (OH-) = 2.4 x 10-13 M



1
Expert's answer
2021-06-14T13:35:21-0400

we know that

"\star""[OH-][H3O+]=1 \times10^{-14}"

"\star P^H=-log[H^+]"

"\star P^{OH}=-log[OH^-]"


1)

"[H_3O^+]" =3.7 x 10-10 M

corresponding [OH-]=2.7 x10-5

"P^H=-log[H^+]=-log[3.7\times10^{-10}]"

=9.44 (>7)

thus, basic.


2)[OH-] = 2.4 x 10-13 M

corresponding [H+]=4.1 x 10-2

"P^H=-log[H^+]=-log[4.1\times10^{-2}]"

=1.387 (<7)

thus ,acidic